Please help me solve this problem
Find the constant k,if the line 3x-4y+5=0 is parallel to kx+6y=-6

Answer by stanbon(75887)

Please help me solve this problem
Find the constant k,if the line 3x-4y+5=0 is parallel to kx+6y=-6
----
Write each equation in the slope intercept form:
y = (3x-5)/4
slope = 3/4
---------------------
y = (-kx-6)/6
slope = -k/6
---------------------
Solve for "k":
-k/6 = 3/4
---
-k = 18/4
---
k = -9/2
================
Cheers,
Stan H.
================